on an average a box containing 10 articles is likely to have 2 defectives. if we consider the consignment of 100 boxes, how many of them are expected to three or less defectives ?
1
Expert's answer
2020-11-01T14:40:43-0500

Binomial distribution

"P=\frac{2}{10}=0.2, n=10"since average "np=2" for 10 articles.

The probability of a box having three or less defectives is"P(\le3)"

"P(\le3)=P(0)+P(1)+P(2)+P(3)"

"=\sum_{x=0}^3\binom{10}{x}P^x(1-P)^{10-x}, x=0,1,2...10"

=0.879

Thus, in a consignment of 100 boxes, "0.879\u00d7100=87.9" which is approximately 88 boxes are expected to have three or less defective articles.